# “6 choose 4”: A framework to understand and facilitate discussion of strategies for overall survival safety monitoring

### v1

@@ Abstract
Advances in anticancer therapies have significantly contributed to declining death rates in certain disease and clinical settings. However, they have also made it difficult to power a clinical trial in these settings with overall survival (OS) as the primary efficacy endpoint. A number of statistical approaches have therefore been recently proposed for the pre-specified analysis of OS as a safety endpoint (Shan, 2023; Fleming et al., 2024; Rodriguez et al., 2024). In this paper, we provide a simple, unifying framework that includes the aforementioned approaches—and a couple others—as special cases. By highlighting each approach’s focus, priority, tolerance for risk, and strengths or challenges for practical implementation, this framework can help to facilitate discussions between stakeholders on “fit-for-purpose OS data collection and assessment of harm” (American Association for Cancer Research, 2023). We apply this framework to a real clinical trial in large B-cell lymphoma to illustrate its application and value. Several recommendations and open questions are also raised.
@@ 1 Introduction
With the continued advancement of anticancer therapies, there are now many disease and clinical contexts in which patients’ lives are extended and multiple treatment options are available to patients following progression. While these successes should be celebrated, they have also made it challenging for clinical trials to have adequate statistical power to detect statistically significant improvements in overall survival (OS)—the gold standard endpoint in oncology (Merino et al., 2024). In July 2023, a workshop titled “Overall Survival in Oncology Clinical Trials” was held jointly by the U.S. Food and Drug Administration (FDA), American Association for Cancer Research (AACR), and American Statistical Association (ASA) to discuss challenges with timely assessment of OS for novel therapies (American Association for Cancer Research, 2023). The workshop cited lack of OS data, lack of plans for further OS data collection, and lack of pre-specified OS analyses as emerging challenges for regulatory authorities to interpret OS data in trials where OS is not a primary or key secondary endpoint. A call-to-action was made to “encourage thoughtful and comprehensive planning during trial design for fit-for-purpose OS data collection and assessment of harm given the disease and clinical context”.
Following that call-to-action, several statistical approaches have been proposed to date for the pre-specified analysis of OS as a safety endpoint. Fleming et al. (2024) proposed a monitoring guideline inspired by ones used by the FDA in cardiovascular safety trials. Rodriguez et al. (2024) suggested using confidence intervals to decide whether or not there is sufficient evidence to rule out harm. Shan (2023) adopted sample size calculation practices for non-inferiority trials to achieve desirable trial operating characteristics.
@@
The existing approaches by Fleming et al. (2024), Rodriguez et al. (2024), and Shan (2023) may seem similar at first glance. For example, all three approaches consider similar sets of parameters, e.g., definition of harm measured in terms of the OS hazard ratio and type I error. However, upon closer look, it can be unclear even to the experienced statistician how these approaches relate to each other, how they differ, which approach one should use in practice, and why.
@@
This paper proposes a framework for OS safety analysis that includes as special cases Fleming et al. (2024), Rodriguez et al. (2024), and Shan (2023). Importantly, this framework explains what flexibility and restrictions exist for any approach analyzing OS as a safety endpoint under the traditional framework of hypothesis testing, and clarifies for each approach their focus, priority, and tolerance for risk over the course of a trial. By increasing understanding of different approaches and the relationship between them, we believe this framework can help facilitate “thoughtful planning” by sponsors and improve communication between stakeholders in order to achieve alignment on clinical trial design and analysis.
@@
The rest of the paper is organized as follows. In Section 2, we present our framework for OS safety monitoring and provide five examples of approaches that fall under this framework, including Fleming et al. (2024), Rodriguez et al. (2024), and Shan (2023). In Section 3, we illustrate and compare the five approaches by applying them to POLARIX, a real clinical trial studying patients with untreated large B-cell lymphoma. Section 4 concludes with a brief summary, recommendations for practice, and open questions for further deliberation or research.
@@ 2 Proposed framework
where $\Phi(\cdot)$ denotes the cumulative distribution function of the standard normal distribution. These equations assume that the observed log hazard ratio $\log(\hat{\theta})$ follows a normal distribution with variance $1/(\pi(1-\pi)d)$ and mean $\log(\theta_{0})$ or $\log(\theta_{1})$ (Schoenfeld, 1981).
Given six parameters and Equations (I)-(II) characterizing the relationship between them, our proposed framework is as follows: at each analysis time, users input or “choose” 4 of the 6 parameters (not all from the same equation) and solve for the remaining 2. This one-sentence framework communicates clearly what flexibility and restrictions exist in any OS monitoring strategy. It highlights a strategy’s focus and associated risks (e.g., choosing $\beta$ and solving for $\alpha$ at an early interim analysis implies a focus on having the trial continue in case the experimental therapy is effective, while understanding the risk of a false positive). And as we shall illustrate, it includes as special cases Fleming et al. (2024), Rodriguez et al. (2024), Shan (2023), and other approaches for OS safety monitoring.
Before proceeding, there are three important points that deserve clarification. First, $\alpha$ and $\beta$ here refer to error rates associated with a single test as opposed to a group-sequential test and/or multiple endpoints. More specifically, they refer to marginal type I and type II error rates for hypothesis testing of OS safety at interim and final analyses. They should not be confused with overall error rates across all analysis times. (Note that Fleming et al. refer to $\alpha$ and $\beta$ as false positive and false negative error rates. We prefer the terms marginal type I and type II error rates to more clearly distinguish them from overall error rates.) They should also not be confused with error rates for primary and key secondary efficacy analyses, which may or may not include OS. Whether OS should be formally tested as a safety endpoint and how to do so in relation to other formally tested endpoints is context specific and deserves careful consideration in practice. However, a thorough discussion of this topic is beyond the scope of this paper.
Second, the proposed framework should not be misunderstood as suggesting a single one-time calculation. Just because a user inputs certain parameters does not mean they should accept whatever the other parameters end up being. All six parameters are important and an iterative process is likely needed to arrive at a configuration that various stakeholders can agree on. The intention of our framework is not to encourage focus on some parameters while ignoring others, but rather to elucidate and encourage stakeholder discussion of the relative priorities and trade-offs between the parameters. Our framework may also help clarify how certain parameters are more informed/guided by data or additional (e.g., clinical or scientific) considerations compared to other parameters.
@@
Fleming et al.’s OS monitoring guideline can be understood under the proposed framework as choosing $(\theta_{0},\theta_{1},\beta,d)$ at interim analyses (IAs) and $(\theta_{0},\theta_{1},\alpha,d)$ at the final analysis (FA). As such, their approach places more emphasis on trial continuation while evidence is still accumulating, which could make sense in trials with limited OS events at early IAs. This is later counter-balanced by a focus on ruling out OS detriment with reasonably low type I error rate $\alpha$ when more OS events have accumulated.
While representing Fleming et al.’s guideline in such a way makes it clearer and simpler to understand, it is worth reiterating that users should avoid strictly interpreting parameters as being either an input or output. Beyond the mechanistic calculations implied by Equations (I)-(II), all parameters are important at each analysis time, whether it’s $\alpha$ to protect patients from detriment, or $\beta$ to ensure that trials of effective drugs are not prematurely terminated, etc.. Therefore, an iterative process is needed to find a “sweet spot” that balances the various parameter values, and we encourage users to think of Fleming et al. and all other approaches under this framework as having different starting lines, different implications with respect to implementation, and (perhaps slightly) different focuses and tolerances for risk.
@@
Rodriguez et al. proposed to calculate the upper bound of the $100\times(1-2\alpha)\%$ confidence interval (CI) for the estimated OS HR, and to conclude low probability of substantial detriment if the upper bound is less than $\theta_{0}$ or high probability of substantial detriment otherwise. This approach is equivalent to choosing $(\theta_{0},\theta_{1},\alpha,d)$ at each analysis time. It intuitively uses CIs to ask, “Can OS detriment be ruled out?”, but risks stopping trials with effective experimental treatments early due to a decreased emphasis on $\beta$. (Note that “choosing” $\theta_{1}$ here does not directly follow from the use of CIs. Rather, it follows from Rodriguez et al.’s consideration of trials which evaluate for OS safety early and then for OS efficacy later when OS is sufficiently powered against a certain alternative $\theta_{1}$.)
Special case #3: Shan (2023)
Shan (2023) proposed to choose $(\theta_{0},\theta_{1},\alpha,\beta)$ and to solve for $(d,\theta^{*})$ in trials with a single readout. This approach is akin to traditional sample size calculation in non-inferiority trials. It can be generalized to trials with multiple readouts by choosing the same set of four parameters $(\theta_{0},\theta_{1},\alpha,\beta)$ at each analysis time. (Of course, users may choose at each time a different set of parameter values, e.g., $\alpha$ may decrease over time.) However, as mentioned before, timing of the first IA (or first multiple IAs) in trials of indolent cancers is often driven by an intermediate or surrogate endpoint, not OS. In the likely situation that the observed number of deaths differs from the prespecified number $d$, it is unclear from this strategy what modifications would be made and how. For cancers with extremely low mortality rate, $d$ deaths may not even be obtainable within a reasonable timeframe.
@@
Another approach—one that we have seen study teams consider—is to choose $(\theta_{0},\theta_{1},d,\theta^{*})$ at each analysis time, with $\theta^{*}$ being some discrete threshold such as 1.0, 1.1, or 1.2. There may be reasons for choosing $\theta^{*}$. Exceeding $\theta^{*}=1.0$, in particular, implies that OS detriment was observed in the trial overall. The OS HR point estimate is also a clinically important result that can have a strong influence on the public’s perception of a drug, providing another reason why sponsors may wish to apply a specific threshold.
However, it is unclear by choosing $(\theta_{0},\theta_{1},d,\theta^{*})$ what risks one is willing to tolerate and how uncertainty due to small number of events is taken into consideration, if at all. Would $\theta^{*}$ need to be changed if a different number of deaths than $d$ is observed? If so, how? $\theta^{*}=1.0$ is also not appropriate if $\theta_{1}$ is close to 1.0, since the marginal type II error rate will be approximately 50% regardless of how many OS events there are.
@@
Our final example is to choose $(\theta_{1},\alpha,\beta,d)$ at IAs and $(\theta_{0},\theta_{1},\alpha,d)$ at the FA. This approach is based on monitoring guidelines that the FDA has used in trials of type 2 diabetes mellitus (T2D) to evaluate cardiovascular risk (FDA Center for Drug Evaluation and Research, 2008), guidelines which also served to inspire approach #1 by Fleming et al. (2024). The key difference between approach #5 and #1 is that, while approach #1 addresses the question “With what certainty can we rule out a particular level of OS detriment?”, approach #5 places more emphasis on type I error control at IAs and addresses the question “What OS detriment can we rule out with a high level of certainty?”.
However, it is worth noting that both approaches lead to the same calculated thresholds $\theta^{*}$, because the thresholds are uniquely determined by the parameters $\theta_{1}$, $\beta$, and $d$. So in fact users do not need to choose between the two. Rather, they can simultaneously use both approaches to understand uncertainty and to evaluate benefit-risk. This can be especially useful when a clinically relevant $\theta_{0}$ has been defined a priori, but few deaths have been observed to enable robust inference with respect to $\theta_{0}$.
@@ 3 Case study
Fleming et al. (2024) illustrated their OS monitoring guideline by retrospectively applying it to POLARIX (Strategy 1, Table 1). Under our proposed framework, their guideline is equivalent to choosing $(d,\theta_{0},\theta_{1},\beta)=(89,1.30,0.80,0.1)$ at IA1, $(d,\theta_{0},\theta_{1},\beta)=(131,1.30,0.80,0.1)$ at IA2, and $(d,\theta_{0},\theta_{1},\alpha)=(178,1.30,0.80,0.025)$ at FA. For the rationale behind some of these parameters, we refer the reader to Fleming et al. (2024). We now illustrate and compare the remaining four approaches described in Section 2 by choosing similar parameter values wherever the set of input parameters overlap with Fleming et al.’s, and choosing conventional/intuitive values wherever they do not.
Table 1: Various OS monitoring strategies, applied to the POLARIX trial and depicted under the proposed framework. For each strategy, the four input or “chosen” parameters at a given stage are indicated in bold. Overall operating characteristics if safety thresholds are strictly adhered to are also provided. Strategy 1 is identical to the case study presented by Fleming et al. (2024), minus one potential interim analysis at 60 deaths for simplicity.
@@
All three 0.018a 0.829b
2A. Rodriguez et IA1 89 1.30 0.80 0.858 0.025 0.629
al. (2024) IA2 131 1.30 0.80 0.923 0.025 0.793
@@
All three 0.007a 0.600b
2B. Rodriguez et IA1 89 1.30 0.80 0.991 0.100 0.843
al. (2024) IA2 131 1.30 0.80 0.975 0.050 0.872
FA 178 1.30 0.80 0.969 0.025 0.900
All three 0.015a 0.783b
3. Shan (2023) IA1 111 1.30 0.80 1.020 0.100 0.900
IA2 145 1.30 0.80 0.990 0.050 0.900
FA 178 1.30 0.80 0.969 0.025 0.900
All three 0.018a 0.841b
@@
guidance in IA2 131 1.41 0.80 1.001 0.025 0.900
T2D (2008) FA 178 1.30 0.80 0.969 0.025 0.900
All three 0.018c 0.829b
aProbability of meeting all three thresholds (IA1, IA2, and FA) under $H_{0}$.
bProbability of meeting all three thresholds (IA1, IA2, and FA) under $H_{1}$.
cProbability of meeting all three thresholds (IA1, IA2, and FA) under HR = 1.30, the HR under $H_{0}$ at FA.
Our first adaptation of Rodriguez et al. (2024) is based on the authors’ suggestion to use OS HR 95% confidence intervals to rule out OS detriment at all analysis times. This is equivalent to specifying $\alpha=0.025$ throughout (Strategy 2.A, Table 1). By emphasizing stringent control of type I error early, it can be seen that the probability of ruling out OS detriment when Pola-R-CHP is effective is substantially lower compared to Strategy 1 (0.629 with 89 deaths and 0.793 with 131 deaths, compared to 0.90 at both IAs).
Rodriguez et al. also suggested that one could relax the type I error rate from 0.025, although they did not provide guidance on how to actually choose $\alpha$. Here, we set $\alpha$ to be 0.10 at IA1, 0.05 at IA2, and 0.025 at FA—levels typical of drug development and which become more stringent as event size increases (Strategy 2.B, Table 2). Doing so results in slightly lower or similar probabilities of meeting the decision making threshold under $H_{1}$ than Strategy 1 (0.843, 0.872, and 0.900).
If there is a desire to control both type I and type II error rates, then one approach could be to use Shan’s to calculate the required event size. For example, Strategy 3 in Table 1 shares the same type I error rates as Strategy 2.B (0.10, 0.05, and 0.025) and the same type II error rates as Strategy 1 (0.10). However, the actual timing of IA in POLARIX was driven by PFS, not OS. Indeed, the trial conducted its analysis when 241 PFS events were observed. If instead the timing was driven by OS, then the readout could have been substantially delayed, with an increased risk of overpowering PFS or sitting on PFS results for many months.
@@
Strategy 4 considers the use of $\theta^{*}=$ 1.1, 1.0, and 1.0 as the HR thresholds for decision making over time. We chose $\theta^{*}=1.1$ for IA1 based on our observation that study teams have a tendency to be conservative early on in a trial when there is little data. Specific to this case study, Strategy 4 can be seen to have similar operating characteristics as Strategy 1. However, the slightly larger type I error rates of 0.215 and 0.040 at IA1 and FA might be a safety concern. More generally, this strategy doesn’t support a transparent discussion on the various risks, since it focuses on the point estimate—which is problematic from a statistical perspective when there are few deaths.
@@
Strategy 5 calculates what $\theta_{0}$ can be ruled out at the IAs with $\alpha=0.025$ and $\beta=0.1$. We see in this case that $\theta_{0}=1.59$ and 1.41 given 89 and 131 deaths, respectively. As noted in Section 2, Strategies 1 and 5 result in the same calculated thresholds $\theta^{*}$. Therefore, a user can use both strategies at an IA to understand with what level of certainty they can rule out $\theta_{0}=1.30$ ($\alpha=0.157$ at IA1 and $\alpha=0.067$ at IA2) and what OS detriment they can rule out with high certainty $\alpha=0.025$ ($\theta_{0}=1.59$ at IA1 and $\theta_{0}=1.41$ at IA2). Using Strategy 5 alone may be problematic, as it doesn’t allow a transparent discussion on the risks associated with a clinically relevant target, and may in fact claim a low risk but for a clinically irrelevant question.
@@
Besides the marginal risks involved at each stage in a trial, stakeholders should also consider the overall risk of the trial should the decision making thresholds $\theta^{*}$ be strictly followed. Table 1 calculates for each strategy the probability of meeting all three thresholds (i.e., probability of consistently being deemed safe) under $H_{0}$ and under $H_{1}$. It can be seen that all strategies have a high probability $(>97\%)$ of being flagged as unsafe at least once (i.e., probability of failing to meet a threshold at least once) during trial conduct if the drug is harmful. Meanwhile, four of the strategies have a $>$80% probability of consistently being deemed safe (i.e., probability of meeting all three thresholds) if the drug is effective, the exceptions being Strategies 2A (60%) and 2B (78.3%).

### v2

@@ Abstract
Advances in anticancer therapies have contributed significantly to decreasing death rates in certain disease and clinical settings. However, they have also made it difficult to power a clinical trial in these settings with overall survival (OS) as the primary efficacy endpoint. Therefore, two approaches have recently been proposed for the pre-specified analysis of OS as a safety endpoint (Fleming et al., 2024; Rodriguez et al., 2024). We provide in this paper a simple unifying framework that includes the aforementioned approaches—and a few others—as special cases. By highlighting each approach’s focus, priority, tolerance for risk, and strengths or challenges for practical implementation, this framework can help facilitate discussions between stakeholders on “fit-for-purpose OS data collection and assessment of harm” (American Association for Cancer Research, 2023). We apply this framework to a real clinical trial in large B-cell lymphoma to illustrate its application and value. Several recommendations and open questions are also raised.
@@ 1 Introduction
With the continued advancement of anticancer therapies, there are now many disease and clinical contexts in which patients’ lives are extended and multiple treatment options are available to patients following progression. Although these successes should be celebrated, they have also made it challenging for clinical trials to have adequate power to detect statistically significant improvements in overall survival (OS)—the gold standard endpoint in oncology (Merino et al., 2024). In July 2023, a workshop titled “Overall Survival in Oncology Clinical Trials” was held jointly by the U.S. Food and Drug Administration (FDA), American Association for Cancer Research (AACR), and American Statistical Association (ASA) to discuss challenges with the timely evaluation of OS for novel therapies (American Association for Cancer Research, 2023). The workshop cited lack of OS data, lack of plans for further OS data collection, and lack of pre-specified OS analyses as emerging challenges for regulatory authorities to interpret OS data in trials where OS is not a primary or key secondary endpoint. A call to action was made to “encourage thoughtful and comprehensive planning during trial design for fit-for-purpose OS data collection and assessment of harm given the disease and clinical context”.
@@
Following that call-to-action, two statistical approaches were proposed for the pre-specified analysis of OS as a safety endpoint. Fleming et al. (2024) proposed a monitoring guideline inspired by ones used by the FDA in cardiovascular safety trials. Rodriguez et al. (2024) suggested using confidence intervals to decide whether or not there is sufficient evidence to rule out harm. These approaches are similar to a certain degree. For example, both consider similar sets of parameters: definition of harm measured in terms of the OS hazard ratio, type I error, etc.. However, upon closer look, it can be unclear even to the experienced statistician how these approaches relate to each other exactly, how they differ, which approach one should use in practice, and why.
@@
This paper proposes a framework for OS safety analysis that includes as special cases Fleming et al. (2024) and Rodriguez et al. (2024). Importantly, this framework explains what flexibility and restrictions exist for any approach analyzing OS as a safety endpoint under the traditional framework of hypothesis testing, and clarifies for each approach their focus, priority, and tolerance for risk over the course of a trial. By increasing understanding of different approaches and the relationship between them, this framework can help facilitate “thoughtful planning” by sponsors and improve communication between stakeholders in order to achieve alignment on clinical trial design and analysis.
@@
The rest of the paper is organized as follows. In Section 2, we present our framework for OS safety monitoring and provide five examples of approaches that fall under this framework, including Fleming et al. (2024) and Rodriguez et al. (2024). In Section 3, we illustrate and compare the five approaches by applying them to POLARIX, a real clinical trial studying patients with untreated large B-cell lymphoma. Section 4 concludes with recommendations for practice and open questions for further deliberation or research.
@@ 2 Proposed framework
[…]
Given six parameters and Equations (I)-(II) characterizing the relationship between them, our proposed framework is as follows: at each analysis time, users input or “choose” 4 of the 6 parameters (not all from the same equation) and solve for the remaining 2. This one-sentence framework communicates clearly what flexibility and restrictions exist in any OS monitoring strategy. It highlights a strategy’s focus and associated risks; for example, choosing $\beta$ and solving for $\alpha$ at an early interim analysis implies a focus on having the trial continue in case the experimental therapy is effective, while understanding the risk of a false positive. And as we shall illustrate, it includes as special cases Fleming et al. (2024), Rodriguez et al. (2024), and other approaches for OS safety monitoring.
Before proceeding, there are three important points that deserve clarification. First, $\alpha$ and $\beta$ here refer to error rates associated with a single test as opposed to a group-sequential test and/or multiple endpoints. More specifically, they refer to marginal type I and type II error rates for hypothesis testing of OS safety at interim and final analyses. They should not be confused with overall error rates across all analysis times. (Note that Fleming et al. refer to $\alpha$ and $\beta$ as false positive and false negative error rates. We prefer the terms marginal type I and type II error rates to more clearly distinguish them from overall error rates.) They should also not be confused with error rates for primary and secondary endpoints with a formal testing plan. Whether OS should be formally tested as a safety endpoint and how to do so in relation to other formally tested endpoints is context specific and deserves careful consideration in practice. However, a thorough discussion of this topic is beyond the scope of this paper.
Second, the proposed framework should not be misunderstood as suggesting a single one-time calculation. Just because a user inputs certain parameters does not mean they should accept whatever the other parameters end up being. All six parameters are important and an iterative process is likely needed to arrive at a configuration that various stakeholders can agree on. The intention of our framework is not to encourage focus on some parameters while ignoring others, but rather to elucidate and encourage stakeholder discussion of the relative priorities and trade-offs between the parameters. Our framework may also help clarify how certain parameters are more informed by data or clinical/scientific considerations compared to other parameters.
@@
Fleming et al.’s OS monitoring guideline can be understood under the proposed framework as choosing $(\theta_{0},\theta_{1},\beta,d)$ at interim analyses (IAs) and $(\theta_{0},\theta_{1},\alpha,d)$ at the final analysis (FA). Their approach places more emphasis on trial continuation while evidence is still accumulating, which could make sense in trials with limited OS events at early IAs. This is later counter-balanced by a focus on ruling out OS detriment with reasonably low type I error rate $\alpha$ when more OS events have accumulated.
While representing Fleming et al.’s guideline in such a way makes it clear and simple to understand, it is worth reiterating that users should avoid strictly interpreting parameters as being either an input or output. Beyond the mechanistic calculations implied by Equations (I)-(II), all parameters are important at each analysis time, whether it’s $\alpha$ to protect patients from harm, or $\beta$ to ensure that trials of effective drugs are not prematurely terminated. Therefore, an iterative process is needed to find a “sweet spot” that balances the various parameter values, and we encourage users to think of Fleming et al. and all other approaches under this framework as having different starting lines, different implications with respect to implementation, and (perhaps slightly) different focuses and tolerances for risk.
@@
Rodriguez et al. proposed to calculate the $100\times(1-2\alpha)\%$ confidence interval (CI) for the estimated OS HR, at a time when OS data is relatively mature. Failure to exclude $\theta_{0}$ with the upper confidence bound is then seen as an indication of OS detriment. Their approach is equivalent to choosing a single $(\theta_{0},\theta_{1},\alpha,\beta)$ and solving for $(d,\theta^{*})$ (Shan, 2023). In other words, like traditional hypothesis testing and sample size calculation, Rodriguez et al. use a CI for decision making precisely at a time when a desirable set of operating characteristics can be achieved.
Rodriguez et al. can be generalized to multiple readouts by choosing the same set of four parameters $(\theta_{0},\theta_{1},\alpha,\beta)$ at each analysis time. (Users will of course have to choose at each time a different set of parameter values, e.g., start with a larger $\alpha$ then decrease it over time.) However, in trials with very low OS event rate, it may be difficult if not impossible to reach the required number of events $d$ within a feasible time frame. In addition, timing of the first IA (or first multiple IAs) in trials of indolent cancers is often driven by an intermediate or surrogate endpoint, not OS. There may not be enough OS data at this time to thoroughly evaluate for potential harm. Even if there is, the observed number of deaths will likely differ from the prespecified number $d$, and it is unclear from this strategy what modifications will be made to address this difference.
Special case #3: Safety monitoring with standard confidence intervals
An alternative approach to Rodriguez et al. (2024) that one might consider is to monitor OS safety using $100\times(1-2\alpha)\%$ CIs again but with flexibility for number of events, i.e., by choosing $(\theta_{0},\theta_{1},\alpha,d)$ at each analysis time. (The choice between $\theta_{1}$ or $\beta$ is flexible and depends on the risk one would like to quantify.)
An example of this approach is an independent data monitoring committee (IDMC) who may not have been provided with detailed instructions for OS safety monitoring and is continually using 95% CIs for evaluation. As commented by Dr. Mikkael Sekeres at the FDA-AACR-ASA workshop, “It is surprising how little guidance you’re often given when you’re on an IDMC… IDMC should be given good guidance on how to consider overall survival… We think that IDMC should be given specific tasks to look at overall survival at different stages of a trial and also statistical guidance on the confidence with which you can declare that a drug is either too harmful or is safe in moving forward” (American Association for Cancer Research, 2023).
However, it is challenging for 95% CIs to rule out harm early on in a trial when OS data is immature and CIs are wide. Expecting them to do so could result in early termination of an effective treatment. Other standard confidence intervals (e.g., 90% or 80%) may be more appropriate for safety monitoring, but specifying $\alpha$ requires careful planning. In particular, one should consider whether or not to change $\alpha$ if a different number of deaths than $d$ is observed, and if so, how.
@@
Another approach—one that we have seen study teams consider—is to choose $(\theta_{0},\theta_{1},d,\theta^{*})$ at each analysis time, with $\theta^{*}$ being some discrete threshold such as 1.0, 1.1, or 1.2. There are pragmatic reasons for choosing $\theta^{*}$. For one, a discrete threshold is easy to communicate, understand, and follow. This could be beneficial for an IDMC that is tasked with assessing OS detriment. In addition, the OS HR point estimate is a clinically important result that can have a strong influence on the public’s perception of a drug, so there could be a desire to cap it. Exceeding $\theta^{*}=1.0$, in particular, implies that OS detriment was observed in the trial overall.
However, it is unclear by choosing $(\theta_{0},\theta_{1},d,\theta^{*})$ what risks one is willing to tolerate and how uncertainty due to small number of events is taken into consideration, if at all. Similar to approach #3, one should consider whether or not $\theta^{*}$ needs to be changed if a different number of deaths than $d$ is observed. $\theta^{*}=1.0$ is also not appropriate if $\theta_{1}$ is close to 1.0, since $\beta$ will be approximately 0.50 regardless of how many OS events there are.
@@
Our final example is to choose $(\theta_{1},\alpha,\beta,d)$ at IAs and $(\theta_{0},\theta_{1},\alpha,d)$ at the FA. This approach is based on monitoring guidelines that the FDA has used in trials of type 2 diabetes mellitus (T2D) to evaluate cardiovascular risk (FDA Center for Drug Evaluation and Research, 2008), guidelines which also served to inspire approach #1 by Fleming et al.. The key difference between approach #5 and #1 is that, while approach #1 addresses the question “With what certainty can we rule out a particular level of OS detriment?”, approach #5 places more emphasis on type I error control at IAs and addresses the question “What OS detriment can we rule out with a high level of certainty?”.
It is worth noting though that both approaches #1 and #5 lead to the same calculated thresholds $\theta^{*}$, because the thresholds are uniquely determined by the parameters $\theta_{1}$, $\beta$, and $d$. Thus, users do not actually need to choose between the two. Rather, they can simultaneously use both approaches to understand uncertainty and to evaluate benefit-risk. This can be especially useful when a clinically relevant $\theta_{0}$ has been defined a priori, but few deaths have been observed to enable robust inference with respect to $\theta_{0}$.
@@ 3 Case study
Fleming et al. (2024) illustrated their OS monitoring guideline by retrospectively applying it to POLARIX (Strategy 1, Table 1). Under our proposed framework, their guideline is equivalent to choosing $(d,\theta_{0},\theta_{1},\beta)=(89,1.30,0.80,0.1)$ at IA1, $(d,\theta_{0},\theta_{1},\beta)=(110,1.30,0.80,0.1)$ at IA2, $(d,\theta_{0},\theta_{1},\beta)=(131,1.30,0.80,0.1)$ at IA3, and $(d,\theta_{0},\theta_{1},\alpha)=(178,1.30,0.80,0.025)$ at FA. For the rationale behind some of these parameters, we refer the reader to Fleming et al. (2024). We now illustrate and compare the remaining four approaches described in Section 2 by choosing similar parameter values wherever the set of input parameters overlap with Fleming et al.’s, and choosing conventional/intuitive values wherever they do not.
Table 1: Various OS monitoring strategies applied to the POLARIX trial and depicted under the proposed framework. For each strategy, the four input or “chosen” parameters at a given stage are indicated in bold. Overall operating characteristics if safety thresholds are strictly adhered to are also provided. The threshold $\theta^{*}$ is considered “met” if the observed OS log hazard ratio falls below it, indicating that the experimental treatment is safe, or at least there is insufficient evidence to claim OS detriment. Strategy 1 is identical to the case study presented by Fleming et al. (2024), minus one potential interim analysis at 60 deaths for simplicity.
@@
Overall 0.017a 0.819b
2. Rodriguez et IA 145 1.30 0.80 0.990 0.050 0.900
al. (2024) FA 178 1.30 0.80 0.969 0.025 0.900
Overall 0.020a 0.869b
3A. Standard IA1 89 1.30 0.80 0.858 0.025 0.629
CIs IA2 110 1.30 0.80 0.895 0.025 0.721
IA3 131 1.30 0.80 0.923 0.025 0.793
@@
Overall 0.006a 0.584b
@@
guidance in IA2 110 1.48 0.80 1.021 0.025 0.900
T2D (2008) IA3 131 1.41 0.80 1.001 0.025 0.900
FA 178 1.30 0.80 0.969 0.025 0.900
Overall 0.017c 0.819b
aProbability of meeting all thresholds (interim and final) under $H_{0}$.
bProbability of meeting all thresholds (interim and final) under $H_{1}$.
cProbability of meeting all thresholds (interim and final) under HR = 1.30, the HR under $H_{0}$ at FA.
If there is a desire to control both type I and type II error rates, then one approach is to follow Rodriguez et al. (2024) by choosing $(\theta_{0},\theta_{1},\alpha,\beta)$ over time. Strategy 2 in Table 1 specifies standard type I error rates that decrease over time (0.05 and 0.025) and type II error rate 0.10. The resulting thresholds $\theta^{*}$ are similar to those of Strategy 1 at IA3 and FA. However, the actual timing of IA in POLARIX was driven by PFS, not OS; the trial targeted 228 PFS events, eventually observing 241 PFS events along with 110 OS events. Should POLARIX have assessed for potential OS detriment then, and if so, how? Or should it have waited until 145 and 178 OS events before assessing for potential OS detriment? The latter may seem odd considering that in reality POLARIX formally assessed OS for efficacy at IA under the group sequential framework.
Strategy 3A gives back flexibility on the timing of safety monitoring by continually using 95% CIs to rule out harm. It is equivalent to choosing $(d,\theta_{0},\theta_{1},\alpha=0.025)$ at all times. Expecting 95% CIs to rule out $\theta_{0}$ early on in POLARIX results in very stringent thresholds (0.858 with 89 deaths, 895 with 110 deaths, and 0.923 with 131 deaths) and substantially lower probabilities of claiming OS safety when Pola-R-CHP is in fact effective (0.629, 0.721, and 0.793). It is for this reason that we do not recommend stringent control of type I error early on in a trial, since immature OS data and wide 95% CIs can cause unnecessary alarm and result in the wrongful termination of effective treatments.
Compared to Strategy 3A, Strategy 3B uses 70%, 80%, and 90% CIs to rule out harm early on in POLARIX. It is equivalent to choosing $(d=89,\theta_{0},\theta_{1},\alpha=0.15)$, $(d=110,\theta_{0},\theta_{1},\alpha=0.1)$, $(d=131,\theta_{0},\theta_{1},\alpha=0.05)$, and $(d=178,\theta_{0},\theta_{1},\alpha=0.025)$. Strategy 3B results in more relaxed thresholds and higher probabilities of claiming OS safety when Pola-R-CHP is effective, similar to those of Strategy 1.
@@
Strategy 4 considers the use of $\theta^{*}=$ 1.1, 1.05, 1.0, and 1.0 as the HR thresholds for decision making over time. We chose $\theta^{*}$ of 1.1 for IA1 and 1.05 for IA2 based on our observation that study teams have a tendency to be conservative early on in a trial when there is little data. Specific to this case study, Strategy 4 can be seen to have similar operating characteristics as Strategy 1. However, the slightly larger type I error rates of 0.215 and 0.040 at IA1 and FA might be a safety concern. More generally, this strategy does not support a transparent discussion on the various risks because it focuses on the point estimate, which can be problematic from a statistical perspective when there are few deaths.
@@
Strategy 5 calculates what $\theta_{0}$ can be ruled out at the IAs with $\alpha=0.025$ and $\beta=0.1$. We see in this case that $\theta_{0}=1.59$, 1.48, and 1.41 given 89, 110, and 131 deaths, respectively. As noted in Section 2, Strategies 1 and 5 result in the same calculated thresholds $\theta^{*}$. Therefore, a user can use both strategies at an IA to understand with what level of certainty they can rule out $\theta_{0}=1.30$ ($\alpha=0.157$ at IA1, $\alpha=0.1031$ at IA2, and $\alpha=0.067$ at IA3) and what OS detriment they can rule out with high certainty $\alpha=0.025$ ($\theta_{0}=1.59$ at IA1, $\theta_{0}=1.48$ at IA2, and $\theta_{0}=1.41$ at IA3). Using Strategy 5 alone though may be problematic, as it doesn’t allow a transparent discussion on the risks associated with a clinically relevant target, and may in fact claim a low risk but for a clinically irrelevant question.
@@
Besides the marginal risks involved at each stage in a trial, stakeholders should also consider the overall risk of the trial should the decision making thresholds $\theta^{*}$ be strictly followed. Table 1 calculates for each strategy the probability of consistently being deemed safe (i.e., probability of meeting all thresholds $\theta^{*}$) under $H_{0}$ and under $H_{1}$. It can be seen that all strategies have a high probability $(>97\%)$ of being flagged as unsafe at least once (i.e., failing to meet a threshold at least once) during trial conduct if the drug is harmful. Meanwhile, five of the strategies have a >>80% probability of consistently being deemed safe if the drug is effective, the only exception being Strategy 3A (58%).